# Addressing Gradient Misalignment in Data-Augmented Training for Robust Speech Deepfake Detection

### v1

@@ 1 Introduction
Speech deepfake detection (SDD) models perform well under seen conditions but struggle to generalize to unseen scenarios. They often suffer from significant accuracy drops under new attack types, acoustic conditions, or noise that can hide synthetic speech artifacts [1, 2]. To address this challenge, data augmentation (DA) is a straightforward strategy to increase training variability for SDD. Common DA techniques in speech processing include room impulse response (RIR) [3, 4], environmental noise addition [5], and masking [6, 7]. Based on that, RawBoost [8] is designed for the SDD task by applying signal processing to distort raw waveforms, which enhances robustness to acoustic variations. Recent studies have explored different SDD augmentation strategies. For example, [9] proposes an adversarial attack-inspired method that produces pseudo-fake samples near decision boundaries. CpAug [10] applies a copy–paste strategy to enhance intra-class diversity. Additionally, [11] explores embedding-level augmentation to diversify the feature space and promote more generalizable representations.
@@
While DA improves the robustness of SDD, prior work [4] has shown that DSD models trained with DA still suffer performance degradation when evaluated under new acoustic conditions within the same test set [4]. This suggests that DA alone is insufficient and better results may be achieved by helping the model focus distinguishing bona fide from spoofed audio rather than overfitting to acoustic variations introduced by augmentation. We hypothesize that the sensitivity of SDD models to DA perturbations leads to different loss surface between original and augmented inputs. As a result, gradients backpropagated from these two input types may point in different directions, causing gradient conflicts during joint updates. These conflicts can slow down convergence, and push optimization toward suboptimal solutions. Similar issues have been observed in multi-task learning [12], where gradients from different tasks can be conflicting and harm training [12]. Gradient alignment methods were originally introduced in multi-task training to reduce such conflicts and stabilize optimization [13, 14, 15]. These methods were later extended to domain generalization to enforce gradient consistency across domains [16].
@@
Motivated by these observations, we propose a dual-path data-augmented (DPDA) training framework with gradient alignment for speech deepfake detection (SDD). In this framework, the model is trained using both original and augmented version of each training utterance, and a gradient alignment method is applied when their backpropagated gradients conflict. Although gradient alignment methods has been studied in other contexts such as multi-task learning and domain generalization, our approach is distinct in that it aligns gradients between two inputs derived from the same utterance but modified through DA. This alignment encourages consistent parameter updates and helps the model focus on spoof-related cues rather than overfitting to augmentation artifacts. To the best of our knowledge, this is the first work to investigate gradient conflicts during data-augmented training in the context of SDD. Our study provides the following key findings: (i) gradient conflicts between the original and augmented inputs are linked to differences in their loss landscapes; (ii) such conflicts occur frequently during DPDA training; and (iii) resolving them consistently improves performance across multiple SDD architectures, augmentation strategies, and benchmark datasets over both our DPDA baseline and conventional single-path data-augmented training. Finally, our approach is architecture-agnostic and can be easily integrated with existing models and signal-level data augmentation techniques.
@@ 2.1 The proposed training framework
Figure 1 illustrates our proposed DAPA training framework with gradient alignment. Let denote a original input speech sample and the augmented counterpart generated through an input-level data augmentation method. Both inputs are formed as a mini-batch and passed through a shared SDD model with parameters to compute the loss . In conventional training, the gradients from both inputs are aggregated directly as . However, although and come from the same utterance, their gradients may differ, and direct aggregation can introduce conflicts that hinder optimization.
To investigate this, we compare the training losses and alongside the gradient magnitudes computed from each input path. As shown in Figure 2, there is a clear pattern that the training loss of the augmented input path is higher than the original one, hence lead to a larger gradient norms of augmented input compared to the orignal one. This imbalance can cause the model updates to be dominated by the augmented input. As a result, this dominance may bias the model toward learning augmentation-specific patterns, rather than focusing on spoof-related cues. The issue can become worse when the gradients from and point in different directions, which can introduce conflicts during optimization.
@@
Figure 3 shows the resulting loss surfaces for both the original and augmented input paths. The loss surface of the original input is relatively smooth, with fewer nearby suboptimal points that have lower training loss than the current model parameters. In contrast, the augmented input produces a more complex surface, with several sharp valleys and suboptimal points, suggesting a more complicated optimization region. We identify the top-3 lowest loss points on each surface to approximate the directions in which the model parameters would ideally move to reduce the training loss, as guided by gradient descent. As can be seen, the directions toward corresponding minima points on the original and augmented loss surfaces are not well aligned, indicating a mismatch in their optimization trajectories. These differences in surface geometry and descent direction reinforce our hypothesis that gradient conflicts arise due to the distinct loss landscapes induced by data augmentation.
These findings highlight the need for gradient alignment during training to keep the optimization direction from both inputs consistent and help the model focus on detecting spoofed speech rather than being influenced by input variations.
@@ 3.1 Dataset and metrics
We train and validate our method on the ASVspoof2019 Logical Access (LA) dataset [18], which consists of bona fide and spoofed speech generated by diverse text-to-speech (TTS) and voice conversion (VC) systems. For evaluation, three datasets including ASVspoof2021 DF (21DF) [19], In-the-Wild (ITW) [20] and the Fake-or-Real (FoR) norm-test subset222kaggle.com/datasets/the-fake-or-real-dataset are utilized. The 2021DF dataset includes a wide range of spoofing attacks and simulate real-world conditions including codec compression and channel variability. The ITW and FoR datasets further tests robustness using real-world audio sourced from online media and podcasts. We report the standard Equal Error Rate (EER) as the main metric.
@@ 3.2 Implementation details
We evaluate the proposed method on multiple model architectures, including XLSR-AASIST [21], XLSR-TCM-Conformer [22], and XLSR-Mamba [23], to validate its generalization. All experiments follow the original training settings of each model. However, the dual-path data-augmented training receives both original and augmented inputs, which roughly doubling memory usage because gradients for both need to be stored. To fit the GPU memory during training, we need to reduce the training mini-batch size from 20 (as used in the original setting) to 10, which includes 5 original and 5 augmented samples. For data augmentation, we use the widely adopted RawBoost method [8]. Unlike other works [21, 22, 23, 24] that train separate models with RawBoost configuration 3 and 5 to evaluate on different test sets, we unify all experiments by applying RawBoost configuration 4. This choice saves computational resources while covering all noise types included in both configurations 3 and 5. No augmentation is applied during validation.
@@ 4.4 Experiments with different data augmentation methods
Table 3 demonstrates that our gradient alignment pipeline consistently improves performance of XLST-Conformer-TCM model across the three types of noise conditions: RawBoost, MUSAN Noise [5] and RIR [3], and their combination. In most test sets, adding PCGrad lowers the EER compared to each baseline, showing that our method can generalize well across different noise augmentation strategies. Notably, models trained with RawBoost both with and without PCGrad achieve better results than those augmented with MUSAN Noise and RIR, indicating the effectiveness of RawBoost for this task. Finally, these findings highlight that our proposed approach is robust and effective for enhancing deepfake detection under various data augmentation methods.

### v2

@@ 1 Introduction
Speech deepfake detection (SDD) models perform well under seen conditions but struggle to generalize to unseen scenarios. They often suffer from significant accuracy drops under new attack types, acoustic conditions, or noise that can hide synthetic speech artifacts [9, 12, 14, 21]. To address this challenge, data augmentation (DA) is a straightforward strategy to increase training variability for SDD. Common DA techniques in speech processing include room impulse response (RIR) [7, 16], environmental noise addition [22], and masking [19, 27]. Based on that, RawBoost [23] is designed for the SDD task by applying signal processing to distort raw waveforms, which enhances robustness to acoustic variations. Recent studies have explored different SDD augmentation strategies. For example, [1] proposes an adversarial attack-inspired method that produces pseudo-fake samples near decision boundaries. CpAug [32] applies a copy–paste strategy to enhance intra-class diversity. Additionally, [5] explores embedding-level augmentation to diversify the feature space and promote more generalizable representations.
@@
While DA improves the robustness of SDD, prior work [16] has shown that SDD models trained with DA still suffer from performance degradation when evaluated under new acoustic conditions within the same test set. This suggests that DA alone is insufficient and better results may be achieved by helping the model focus on distinguishing bona fide from spoofed audio rather than overfitting to acoustic variations introduced by augmentation. We hypothesize that the sensitivity of SDD models to DA perturbations leads to different loss surfaces between original and augmented inputs. As a result, gradients backpropagated from these two input types may point in different directions, causing gradient conflicts during joint updates. These conflicts can slow down convergence, and push optimization toward suboptimal solutions. Similar issues have been observed in multi-task learning [2], where gradients from different tasks can be conflicting and harm training. Gradient alignment methods were originally introduced in multi-task training to reduce such conflicts and stabilize optimization [31, 29, 11]. These methods were later extended to domain generalization to enforce gradient consistency across domains [17].
@@
Motivated by these observations, we propose a dual-path data-augmented (DPDA) training framework with gradient alignment for speech deepfake detection (SDD). In this framework, the model is trained using both original and augmented version of each training utterance, and a gradient alignment method is applied when their backpropagated gradients conflict. Although gradient alignment methods have been studied in other contexts such as multi-task learning and domain generalization, our approach is distinct in that it aligns gradients between two inputs derived from the same utterance but modified through DA. This alignment encourages consistent parameter updates and helps the model focus on spoof-related cues rather than overfitting to augmentation artifacts. To the best of our knowledge, this is the first work to investigate gradient conflicts during data-augmented training in the context of SDD. Our study provides the following key findings: (i) gradient conflicts between the original and augmented inputs are linked to differences in their loss landscapes; (ii) such conflicts occur frequently during DPDA training; and (iii) resolving them consistently improves performance across multiple SDD architectures, augmentation strategies, and benchmark datasets over both our DPDA baseline and conventional single-path data-augmented training. Finally, our approach is architecture-agnostic and can be easily integrated with existing models and signal-level data augmentation techniques.
@@ 2.1 The proposed training framework
Figure 1 illustrates our proposed DPDA training framework with gradient alignment. Let denote the original input sample and the augmented counterpart generated through an input-level data augmentation method. Both inputs are formed as a mini-batch and passed through a shared SDD model with parameters to compute the loss . In conventional training, the gradients from both inputs are aggregated directly as . However, although and come from the same utterance, their gradients may differ, and direct aggregation can introduce conflicts that hinder optimization.
To investigate this, we compare the training losses and alongside the gradient magnitudes computed from each input path. As shown in Figure 2, there is a clear pattern that the training loss of the augmented input path is higher than the original one, hence leads to a larger gradient norms of the augmented input compared to the original one. This imbalance can cause the model updates to be dominated by the augmented input. As a result, this dominance may bias the model toward learning augmentation-specific patterns, rather than focusing on spoof-related cues. The issue can become worse when the gradients from and point in different directions, which can introduce conflicts during optimization.
@@
Figure 3 shows loss surfaces for both the original and augmented input paths. The loss surface of the original input is relatively smooth, with fewer nearby suboptimal points that have lower training loss than the current model parameters. In contrast, the augmented input produces a more complex surface, with several sharp valleys and suboptimal points, suggesting a more complicated optimization region. We identify the top-3 lowest loss points on each surface to approximate the directions in which the model parameters would ideally move to reduce the training loss, as guided by gradient descent. As can be seen, the directions toward corresponding minima points on the original and augmented loss surfaces are not well aligned, indicating a mismatch in their optimization trajectories. These differences in surface geometry and descent direction reinforce our hypothesis that gradient conflicts arise due to the distinct loss landscapes induced by data augmentation. These findings highlight the need for gradient alignment during training to keep the optimization direction from both inputs consistent and help the model focus on detecting spoofed speech rather than being influenced by input variations.
@@ 3.1 Dataset and metrics
We train and validate our method on the ASVspoof2019 Logical Access (LA) dataset [25], which consists of bona fide and spoofed speech generated by diverse text-to-speech (TTS) and voice conversion (VC) systems. For evaluation, three datasets including ASVspoof2021 DF (21DF) [15], In-the-Wild (ITW) [18] and the Fake-or-Real (FoR) norm-test subset222kaggle.com/datasets/the-fake-or-real-dataset are utilized. The 2021DF dataset includes a wide range of spoofing attacks and simulates real-world conditions including codec compression and channel variability. The ITW and FoR datasets further test robustness using real-world audio sourced from online media and podcasts. We report the standard Equal Error Rate (EER) as the main metric.
@@ 3.2 Implementation details
We evaluate the proposed method on multiple model architectures, including XLSR-AASIST [24], XLSR-TCM-Conformer [26], and XLSR-Mamba [30], to validate its generalization. All experiments follow the original training settings of each model. However, the dual-path data-augmented training receives both original and augmented inputs, which roughly doubling memory usage, as gradients for both need to be stored. To fit the GPU memory during training, we need to reduce the training mini-batch size from 20 (as used in the original setting) to 10, which includes 5 original and 5 augmented samples. For data augmentation, we use the widely adopted RawBoost method [23]. Unlike other works [24, 26, 30, 20, 8] that train separate models with RawBoost configuration 3 and 5 to evaluate on different test sets, we unify all experiments by applying RawBoost configuration 4. This choice saves computational resources while covering all noise types included in both configurations 3 and 5. No augmentation is applied during validation.
@@ 4.4 Experiments with different data augmentation methods
Table 3 demonstrates that our gradient alignment pipeline consistently improves performance of XLSR-Conformer-TCM model across the three types of noise conditions: RawBoost, MUSAN Noise [22] and RIR [7], and their combination. In most test sets, adding PCGrad lowers the EER compared to each baseline, showing that our method can generalize well across different noise augmentation strategies. Notably, models trained with RawBoost both with and without PCGrad achieve better results than those augmented with MUSAN Noise and RIR, indicating the effectiveness of RawBoost for this task. Finally, these findings highlight that our proposed approach is robust and effective for enhancing deepfake detection under various data augmentation methods.